OpenStudy (anonymous):

x[n]=cos((pi/8)*(n^2)). whether it is periodic or not? what is the period of this?

OpenStudy (kc_kennylau):

\[x[n]=\cos\left(\frac\pi8\times n^2\right)\]

OpenStudy (kc_kennylau):

WolframAlpha says no